Reserve Capacity (RC) is the number of minutes a fully charged battery at 80 degrees will discharge 25 amps until the battery drops below 10.5 volts. An automobile manufacturer wants to compare the mean RC for batteries from two different suppliers. RC data (in minutes) for samples of 50 batteries from each supplier are shown below. Construct a 90% confidence interval for the difference between the mean RC for supplier 1 and the mean RC for supplier 2.

Supplier 1 Supplier 2
79.1832 77.1428
71.5732 83.7728
90.2032 73.9528
76.6432 74.9428
88.1932 77.4628
60.2732 78.9128
78.7832 91.4228
64.4332 83.7528
72.9632 76.2728
97.4032 86.4828
88.7832 75.7528
84.7832 66.6928
78.0432 85.3528
74.4832 72.0528
87.9332 71.6528
81.8432 90.0728
82.2432 62.0128
97.9932 78.7928
67.9032 86.7828
77.3832 71.1628
102.5232 101.7328
74.1632 90.9028
87.6532 88.6628
72.0332 80.4028
70.1332 81.8828
82.8732 83.9328
97.0332 75.0328
88.4332 82.1628
94.6132 90.2728
85.0732 72.2028
81.1932 80.2628
82.1432 79.7728
84.1632 80.7828
70.8732 76.0228
79.6732 88.9028
63.1832 76.1928
103.1632 71.6028
63.7432 102.4228
81.2432 83.1328
97.8432 68.0628
90.1332 81.7928
89.5932 66.0028
82.1332 60.1828
92.1632 86.9928
100.3132 73.0528
77.6032 85.3528
84.1032 88.7128
78.7932 53.8928
76.8932 75.8028
74.4732 71.9228

a) What is the lower limit of the 90% confidence interval? Give your answer to three decimal places.

b) What is the upper limit of the 90% confidence interval? Give your answer to three decimal places.

c) Based on this interval, does the claim that mean RC rating is the same for both suppliers seem reasonable?

a. No because 0 is inside the interval.

b. No because 0 is not inside the interval.

c. Yes because 0 is not inside the interval.

d. Yes because 0 is inside the interval.
Answer #1

Let n and nz be the sizes of sample Supplier 1 and Supplier 2 respectively noso ne so be the mean of suppliex 1 X Š Xi 4109 8

a) Lower limit - (x,-X2) - 2012 S2 S2 ni + hz - 682.18 - 79,25) - 1.64 x 109.9674 90.7491 + So So = 2,93 -1,64 x 2,0035 = 2.9